an electron enters a magnetic field vertically downwards with a velocity v from east.the electron is deflected along?
 

Grade:Select Grade

3 Answers

KAPIL MANDAL
9 years ago
The force acting on them is F= e(v×B). Now, charges of electron and proton have opposite sign. Again from Newtons law, F=ma where a is the acceleration. So, 
   e(v×B)=ma. 
 a=e(v×B)/m
Now, charges of electron and proton have opposite sign, So,
firstly, the direction of them will be opposite.
Secondly, mass m of proton is 1836 times greater than that of the electron. So, a=magnitude of acceleration of proton will be 1836 times smaller. Ok? 
KAPIL MANDAL
9 years ago
It is v×B. So, along (-x)×(-z).
So, it is -y direction, or here, north.
KAPIL MANDAL
9 years ago
I’m extremely sorry. I didn’t considered the negative charge of electron. taking a minus in above answer, the answer becomes south.
Is it ok?